OpenStudy (anonymous):

What is the solution of the system? Use substitution. y = x + 6 y = 2x

OpenStudy (anonymous):

2x = x + 6 x = 6

OpenStudy (anonymous):

A. (6, 12) B. (–12, –6) C. (–6, –12) D. (2, 4)

OpenStudy (anonymous):

y = 2(6) y = 12

OpenStudy (anonymous):

A

OpenStudy (anonymous):

What is the solution of the system? Use substitution. x = –2y + 10 –6y = 6x – 6 z = –3x + 10y

OpenStudy (anonymous):

-6y = 6(-2y + 10) - 6 Solve for why then plug your answer into x = -2y+10. Once you have both x and y plug those numbers into the z equation.

OpenStudy (anonymous):

y* not why lol